OpenStudy (anonymous):

If x is nine less than seven and y is nine less seven , what is x less y ? Please show me the steps.

OpenStudy (nenadmatematika):

x=7-9 so x=-2 y=9-7 so y=2 and x-y=-4

OpenStudy (anonymous):

Thank you so much!

OpenStudy (nenadmatematika):

you're welcome :D